OpenStudy (kanwal32):

how to write the equation of projectile after it has reached its maximum height(suppose at time t1 and t2 there are at same height) without finding the the time taken to reach at it full height

OpenStudy (kanwal32):

@Michele_Laino

OpenStudy (kanwal32):

@Michele_Laino pls help in hurry i want to solve this doubt

OpenStudy (kanwal32):

is there any way

OpenStudy (michele_laino):

in general I solve that type of problem using vectors equations, namely: \[\large \left\{ \begin{gathered} {\mathbf{OP}}\left( t \right) = {\mathbf{O}}{{\mathbf{P}}_{\mathbf{0}}} + {{\mathbf{v}}_{\mathbf{0}}}t + \frac{1}{2}{\mathbf{g}}{t^2} \hfill \\ {\mathbf{v}}\left( t \right) = {{\mathbf{v}}_{\mathbf{0}}} + {\mathbf{g}}t \hfill \\ \end{gathered} \right.\]

OpenStudy (michele_laino):

Then I consider a reference frame and I will rewrite those equations in that reference frame

OpenStudy (michele_laino):

for example, if we pick the subsequent reference frame: 

OpenStudy (michele_laino):

then those vector equation can be rewritten as follows: \[\Large \begin{gathered} \left\{ \begin{gathered} z\left( t \right) = {z_0} + {v_0}\left( {\sin \theta } \right)t - \frac{1}{2}g{t^2} \hfill \\ x\left( t \right) = {x_0} + {v_0}\left( {\cos \theta } \right)t \hfill \\ \end{gathered} \right. \hfill \\ \hfill \\ \left\{ \begin{gathered} {v_z}\left( t \right) = {v_0}\left( {\sin \theta } \right) - gt \hfill \\ {v_x}\left( t \right) = {v_0}\left( {\cos \theta } \right) \hfill \\ \end{gathered} \right. \hfill \\ \end{gathered} \]